# Magnetorheological Characterization of Blood Analogues Seeded with Paramagnetic Particles

### v1

@@ Abstract
Abstract - Magnetic particles can be useful in various medical applications, gaining access to the whole body if deployed in the blood stream. Localised drug delivery, haemorrhage control and cancer treatment are among the applications with potential to become revolutionary therapies. Despite this interest, a magnetorheological characterisation of particle-seeded blood has yet to be achieved. In this work, we evaluate the magnetorheological response of polymeric blood analogues seeded with paramagnetic particles in different concentrations, under the effects of a uniform, density-varying magnetic field. Through steady shear experiments, we encounter the usual magnetically-induced shear thinning response, and oscillatory shear results point toward some significant alterations in the fluids’ microstructure. However, experimental limitations make it difficult to accurately evaluate the oscillatory shear response of such rheologically subtle fluids, limiting both the quality and quantity of achievable information. Despite experimental limitations, our results demonstrate that magnetic fields can induce marked and quantifiable rheological changes in seeded blood analogues. The framework established here provides a foundation for future studies on real blood samples and for the design of magnetically responsive biomedical systems
@@ 1 Introduction
It has long been known that blood has magnetic properties1. Red blood cells (RBCs or erythrocytes) are one the major blood constituents and are the main responsible for its mechanical properties2. RBCs contain haemoglobin, a protein whose iron content makes it paramagnetic when not carrying oxygen (deoxyhaemoglobin)1, 3. When under a magnetic field, dipole interactions arise between RBCs, leading them to aggregate into chains aligned with the field. This can be used to knowingly alter blood’s rheological properties without resourcing to drugs or other possibly harmful means4, which is extremely attractive as blood rheology is connected to a series of paramount conditions and diseases. Heart attack, stroke, aneurysms, hypertension and other, possibly crippling or even fatal conditions are associated with high blood viscosity and other rheological characteristics5, 6. On the other hand, the increase of viscosity may hinder blood flow to aid in controlling haemorrhages in complex injuries or surgery7.
Additionally to the direct alteration of haemorheological properties, magnetic fields can be used in a multitude of medical applications8. Among these, incorporating magnetic particles into blood flow holds significant promise for advanced medical treatments, such as targeted drug delivery, haemorrhage control and cancer treatment through hyperthermia therapies or localised blood/oxygen starvation9, 10, 11, 12, 8. However, despite these potential benefits, a comprehensive magnetorheological characterization of blood loaded with magnetic particles has yet to be achieved. This study aims to address this gap by thoroughly analysing the behaviour of both Newtonian and viscoelastic blood analogues under magnetic fields, while also assessing the influence of factors such as particle type and concentration.
@@ 1.1 Magnetorheology
Magnetorheological (MR) fluids are usually suspensions of magnetisable particles in a non-magnetisable continuous phase. Under a magnetic field, dipole interactions between the particles will force them to aggregate into elongated chains aligned with the field lines, and the induced microstructure leads to significant alterations of the rheological properties13. Typically, magnetised MR fluids present a solid-like behaviour at rest, where a stress threshold must be surpassed to induce fluid flow (yield stress). Considering steady shear, a strong shear-thinning response is usually observed as the shear rate is increased due to the enhanced viscous shearing forces breaking down the microstructure. Regarding the response to oscillatory shear, MR fluids typically exhibit a small linear viscoelastic regime (LVE) with a predominantly elastic response (first harmonic storage modulus larger than the loss modulus, $G^{\prime}_{1}>G^{\prime\prime}_{1}$). Increasing the strain typically leads to a steady decrease of $G^{\prime}_{1}$, while $G^{\prime\prime}_{1}$ reaches a maximum before decreasing more slowly than $G^{\prime}_{1}$14, 15, 16, 17. As such, MR fluids usually display a type III behaviour18, where the overshoot in $G^{\prime\prime}_{1}$ at small strain has been attributed to small rearrangements of the unstable microstructure17, 14, dissipating energy into the continuous phase and increasing $G^{\prime\prime}_{1}$. Further increasing the strain leads to a decrease of $G^{\prime\prime}_{1}$ as significant microstructural changes occur within the oscillatory cycle, particularly the breakdown of the microstructure at large instantaneous strain rates. At small oscillation frequencies, the response is practically frequency-independent, and a set of rheological transitions can be defined with strain increase. At sufficiently small deformations, the response is linearly viscoelastic, and increasing the strain onsets a non-linear viscoelastic behaviour from the slight rearrangements of the microstructure. Further increasing the strain results in large microstructural changes within the oscillatory cycle, which results in a deeply non-linear response until, at sufficiently large deformation, the elastic contribution is very small, and the response becomes viscoplastic. At larger frequencies, a Newtonian response is expected with increasing strain from negligible microstructure re-formation until, at a sufficiently large frequency, a Newtonian response is found independently of the strain14, 17.
@@ 1.2 FT-Chebyshev analysis
The non-linear viscoelastic (NLVE) behaviour of MR fluids has been extensively discussed, but the intricacies of the non-linear response remain obscure in measures of average viscoelasticity ($G^{\prime}_{1}$ and $G^{\prime\prime}_{1}$). To uncover the complexities of non-linear viscoelasticity, (Fourier transform) FT-Chebyshev analysis comes into play. For the sake of clarity, we will not discuss the mathematical foundation of FT-Chebyshev formulation; for an in-depth description, we refer to the work of Ewoldt et al.. Nevertheless, a few concepts are worthy of mentioning here. In a standard oscillatory shear test, a sinusoidal strain is imposed on the sample, $\gamma(t)=\gamma_{0}\,\sin(\omega\,t)$, where $\gamma_{0}$ is the strain amplitude and $\omega$ the oscillation frequency. The strain rate is then given by the time derivative, $\dot{\gamma}(t)=\dot{\gamma}_{0}\,\cos(\omega\,t)$, orthogonal to the strain, where $\dot{\gamma}_{0}=\gamma_{0}\,\omega$ is the strain rate amplitude. In a small amplitude oscillatory shear test (SAOS), the measurement is conducted within the LVE where the stress response is linear, and the usual storage and loss moduli ($G^{\prime}_{1}$ and $G^{\prime\prime}_{1}$) can be used to characterise the material. However, sufficiently increasing the strain into large amplitude oscillatory shear (LAOS) leads to a non-linear behaviour where additional harmonics distort the stress response, which is no longer a simple sinusoid. The stress can then be written as a Fourier series of multiple harmonic contributions, which can be used to identify non-linear behaviour. However, this methodology lacks the physical meaning of the higher-order contributions. At the same time, the resulting first harmonic moduli ($G^{\prime}_{1}$ and $G^{\prime\prime}_{1}$) are inadequate in the non-linear regime and give only a measure of the average viscoelastic behaviour19.
@@
Another way of probing non-linearity is through the elastic and viscous Lissajous (or Lissajous–Bowditch) curves, in which the raw stress is plotted as a function of the strain and strain rate, respectively. A Hookean solid would return a positive diagonal elastic Lissajous curve and a perfectly circular viscous Lissajous curve, whereas a purely viscous material would yield the inverse. A linear viscoelastic response is then given by perfect ellipses in both elastic and viscous Lissajous, and distortions of the ellipse shape characterise a non-linear response, as higher harmonics affect the stress signal. Cho et al. proposed decomposing the raw stress in LAOS tests into elastic and viscous components, which effectively collapse the Lissajous into one-dimensional curves, facilitating the analysis. Additionally, the Lissajous curves can also provide quantitative information. Ewoldt et al. defined the minimum strain modulus as the tangent modulus at minimum strain, $G^{\prime}_{M}=d\sigma/d\gamma|_{\gamma=0}$, and the large strain modulus as the secant modulus at maximum strain, $G^{\prime}_{L}=\sigma/\gamma|_{\gamma=\gamma_{0}}$. Similarly, minimum and large rate dynamic viscosities can be defined as $\eta^{\prime}_{M}=d\sigma/d\dot{\gamma}|_{\dot{\gamma}=0}$ and $\eta^{\prime}_{L}=\sigma/\dot{\gamma}|_{\dot{\gamma}=\dot{\gamma}_{0}}$, respectively. As such, in the non-linear regime, these measures help quantify the (intracycle) non-linear behaviour, while in the linear regime, they reduce to $G^{\prime}_{1}$ and $\eta^{\prime}_{1}=G^{\prime\prime}_{1}/\omega$, respectively.
@@
Ewoldt et al. also proposed to fit Chebyshev polynomials of the first kind to the decoupled stresses to allow an analysis of the harmonic contributions independently of the number of selected harmonics, which is usually arbitrary. The first order elastic and viscous Chebyshev coefficients are related to the widely used first harmonic moduli: $e_{1}=G^{\prime}_{1}$ and $v_{1}=\eta^{\prime}_{1}=G^{\prime\prime}_{1}/\omega$, respectively, which are the sole contributors in the LVE; and the non-linear response can be described through the third order coefficients. A positive third-order elastic Chebyshev coefficient, $e_{3}>0$, relates to a higher elastic stress at maximum strain than the linear prediction (considering $e_{1}$ alone), whereas a negative value, $e_{3}<0$, corresponds to a lesser elastic stress at maximum strain than the linear representation. The same can be gathered from the third-order viscous coefficient regarding the viscous stress to the strain rate. This has led $e_{3}>0$ and $e_{3}<0$ to be respectively associated with intracycle elastic strain stiffening and softening, and $v_{3}>0$ and $v_{3}<0$, to intracycle viscous shear thickening and thinning, which often contradict the average viscoelastic behaviour described by $G^{\prime}_{1}$ and $G^{\prime\prime}_{1}$. This issue has been addressed before21, 22 and an alternative framework has been suggested that allows the interpretation of the third order coefficients within the context of the first harmonic non-linear behaviour22. Because the strain and strain rate are orthogonal by definition, the sign of the third order coefficients may be understood, not as a description of the non-linear response itself, but as an indicator of the deformation input that leads to the average non-linearity. For example, a positive $e_{3}$ relates to a relatively larger elastic stress at maximum strain ($\gamma=\gamma_{0}$) or minimum rate ($\dot{\gamma}=0$); as such, if the material presents elastic softening (that is, a decrease of $G^{\prime}_{1}$), $e_{3}>0$ indicates that the elastic softening is driven by the strain rate. The same can be said of the viscous response; a positive $v_{3}$ with an average viscous thinning material ($G^{\prime\prime}_{1}$ decrease) relates to the thinning non-linearity being driven by the strain, as the viscous stress is relatively lesser at $\gamma=\gamma_{0}$.
@@
FT-Chebyshev analysis is a powerful tool to unveil and quantify the intricacies of a material’s non-linear response. However, there is still a lack of in-depth analysis of non-linear magnetorheological data in the literature. As far as we know, the work of Wang et al. is one of the few studies that probe experimental magnetorheological data through FT-Chebyshev analysis. The authors performed LAOS tests with a magnetorheological grease (a suspension of carbonyl iron particles in a semi-solid grease) under different magnetic field intensities. The authors found the application of the magnetic field transformed the rheological response from type I to type III18, and the LVE was significantly shortened, widening with magnetic field intensity increase but still narrower than for the non-magnetised grease. In the absence of magnetic influence, the MR grease is already highly viscoelastic, displaying elastic softening and viscous thinning through the decrease of $G^{\prime}_{1}$ and $\eta^{\prime}_{1}$, both driven by the strain rate ($G^{\prime}_{M}<G^{\prime}_{L}$ and $e_{3}>0$, $\eta^{\prime}_{L}<\eta^{\prime}_{M}$ and $v_{3}<0$). The magnetic field application led to significant alterations in the non-linear response. The authors report a sign shift of $e_{3}$ to negative values at larger strain (for intermediate fields) and positive $v_{3}$ at small strain, later shifting to $v_{3}<0$. The authors discuss the possibility of complex response arising from interactions between the fibre structure of the continuous phase and the magnetically-induced particle-chain structure.

### v2

@@ Abstract
Abstract - Magnetic particle under external fields can be useful in various medical applications, gaining access to the whole body if deployed in the bloodstream. Localised drug delivery, haemorrhage control, and cancer treatment are among the applications that have the potential to become revolutionary therapies. Despite this interest, a magnetorheological characterisation of particle-seeded blood has yet to be achieved. In this work, we evaluate the magnetorheological response of blood analogues seeded with paramagnetic particles in different concentrations, under the effects of a uniform, density-varying magnetic field. Through steady shear experiments, we encounter the usual magnetically-induced shear thinning response, and oscillatory shear results point toward significant alterations in the fluids’ microstructure. However, experimental limitations make it difficult to accurately evaluate the oscillatory shear response of such rheologically subtle fluids, limiting both the quality and quantity of achievable information. Despite experimental limitations, our results demonstrate that magnetic fields can induce marked and quantifiable rheological changes in seeded blood analogues. The framework established here provides a foundation for future studies on real blood samples and for the design of magnetically responsive biomedical systems.
@@ 1 Introduction
It has long been known that blood has magnetic propertiespauling1936. Red blood cells (RBCs or erythrocytes) are one the major blood constituents and are the main responsible for its mechanical propertiesrobertson2008. RBCs contain haemoglobin, a protein whose iron content makes it paramagnetic when not carrying oxygen (deoxyhaemoglobin)pauling1936, zborowski2003. When under a magnetic field, dipole interactions arise between RBCs, leading them to aggregate into chains aligned with the field. This can be used to knowingly alter blood’s rheological properties without resourcing to drugs or other possibly harmful meanstao2011, which is extremely attractive as blood rheology is connected to a series of paramount conditions and diseases. Heart attack, stroke, aneurysms, hypertension and other, possibly crippling or even fatal conditions are associated with high blood viscosity and other rheological characteristicsdintenfass1963, campo2015. On the other hand, the increase of viscosity may hinder blood flow to aid in controlling haemorrhages in complex injuries or surgerymisra2009.
Blood is, however, a very complex fluid and the effects of magnetic field influence on its microstructure and behaviour are still under investigation. The effects on the viscoelastic properties of blood are relatively undiscussed as most experimental works focus on viscosity, which still return conflicting results. tao2011 report that a short magnetic field pulse along the flow direction (1.3 T for approximately 1 min) induced a significant viscosity decrease (20-30%), lasting a few hours before returning to the original value. yan2024 also observed a significant viscosity decrease with a capillary viscometer under an alternating magnetic field. tao2011 attribute the viscosity decrease to magnetic-induced RBC aggregation into larger, streamlined structures and increased polydispersity, while yan2024 additionally contemplate the magnetically-enhanced rheological properties of the RBCs membranes, which facilitates inter-cell sliding that further decreases resistance to flow. On the other hand, haik2001 observed an increase in viscosity of venous blood (30%) when subject to a strong magnetic field parallel to the flow (10 T), ascribing it to magnetic torque maintaining an angular velocity difference between the RBCs and the plasma, leading to an additional viscous dissipation. The contradicting literature dictates that more research needs to be conducted and, particularly, the details of the experimental setup and properties of the blood samples need to be thoroughly described as the magnetorheological response can be heavily influenced by flow dimensions, magnetic field properties, temperature, RBC concentration and blood oxygenationyamamoto2004, yan2024, fahraeus1931.
In addition to the direct alteration of haemorheological properties, magnetic fields can be used in a multitude of medical applicationsandra2007. Among these, incorporating magnetic particles into blood flow holds significant promise for advanced medical treatments, such as targeted drug delivery, haemorrhage control and cancer treatment through hyperthermia therapies or localised blood/oxygen starvationliu2001, arruebo2007, tekleab2019, andra2007. However, despite these potential benefits, a comprehensive magnetorheological characterization of blood loaded with magnetic particles has yet to be achieved. This study aims to address this gap by thoroughly analysing the behaviour of both Newtonian and viscoelastic blood analogues under magnetic fields, while also assessing the influence of factors such as particle type and concentration.
@@ 1.1 Magnetorheology
Magnetorheological (MR) fluids are usually suspensions of magnetisable particles in a non-magnetisable continuous phase. Under a magnetic field, dipole interactions between the particles will force them to aggregate into elongated chains aligned with the field lines, and the induced microstructure leads to significant alterations of the rheological propertiesginder1998. Typically, magnetised MR fluids present a solid-like behaviour at rest, where a stress threshold must be surpassed to induce fluid flow (yield stress). Considering steady shear, a strong shear-thinning response is usually observed as the shear rate is increased due to the enhanced viscous shearing forces breaking down the microstructure. This competition between viscous and magnetic effects has been widely characterised by the Mason number, $Mn$, and the response to steady shear can be reasonably described by the Casson model, which has been found to better fit experimental data than the Bingham model, particularly through a smoother transition from the yielding to the Newtonian responseberli2012, morillas2019.
Regarding the response to oscillatory shear, MR fluids typically exhibit a small linear viscoelastic regime (LVE) with a predominantly elastic response (storage modulus larger than the loss modulus, $G^{\prime}>G^{\prime\prime}$). Increasing the strain typically leads to a steady decrease of $G^{\prime}$, while $G^{\prime\prime}$ reaches a maximum before decreasing more slowly than $G^{\prime}$li2004, sim2003, vicente2011review, parthasarathy1999. As such, MR fluids usually display a type III behaviourhyun2002, where the overshoot in $G^{\prime\prime}$ at small strain has been attributed to small rearrangements of the unstable microstructureparthasarathy1999, li2004, dissipating energy into the continuous phase and increasing $G^{\prime\prime}$. Further increasing the strain leads to a decrease of $G^{\prime\prime}$ as significant microstructural changes occur within the oscillatory cycle, particularly the breakdown of the microstructure at large instantaneous strain rates. At small oscillation frequencies, the response is practically frequency-independent, and a set of rheological transitions can be defined with strain increase. At sufficiently small deformations, the response is linearly viscoelastic, and increasing the strain onsets a non-linear viscoelastic behaviour from the slight rearrangements of the microstructure. Further increasing the strain results in large microstructural changes within the oscillatory cycle, which results in a deeply non-linear response until, at sufficiently large deformation, the elastic contribution is very small, and the response becomes viscoplastic. At larger frequencies, a Newtonian response is expected with increasing strain from negligible microstructure re-formation until, at a sufficiently large frequency, a Newtonian response is found independently of the strainli2004, parthasarathy1999.
@@ 1.2 FT-Chebyshev analysis
The non-linear viscoelastic (NLVE) behaviour of MR fluids has been extensively discussed, but the intricacies of the non-linear response remain obscure in measures of average viscoelasticity ($G^{\prime}$ and $G^{\prime\prime}$). To uncover the complexities of non-linear viscoelasticity, (Fourier transform) FT-Chebyshev analysis comes into play. For the sake of clarity, we will not discuss the mathematical foundation of FT-Chebyshev formulation; for an in-depth description, we refer to the work of ewoldt2008. Nevertheless, a few concepts are worth mentioning here. In a standard oscillatory shear test, a sinusoidal strain is imposed on the sample, $\gamma(t)=\gamma_{0}\,\sin(\omega\,t)$, where $\gamma_{0}$ is the strain amplitude and $\omega$ the oscillation frequency. The strain rate is then given by the time derivative, $\dot{\gamma}(t)=\dot{\gamma}_{0}\,\cos(\omega\,t)$, orthogonal to the strain, where $\dot{\gamma}_{0}=\gamma_{0}\,\omega$ is the strain rate amplitude. In a small amplitude oscillatory shear test (SAOS) the measurement is conducted within the LVE where the stress response is linear, which can fully described by the first-harmonic storage and loss moduli ($G^{\prime}_{1}$ and $G^{\prime\prime}_{1}$)iiiThe viscoelastic moduli until now represented by $G^{\prime}$ and $G^{\prime\prime}$ referred to these first-harmonic moduli.. However, increasing the strain into the large amplitude oscillatory shear regime (LAOS) leads to a non-linear behaviour where additional harmonics distort the stress response, which is no longer a simple sinusoid. The stress can then be written as a Fourier series of multiple harmonic contributions, which can be used to identify non-linear behaviour. However, this methodology lacks the physical meaning of the higher-order contributions. At the same time, the resulting first-harmonic moduli ($G^{\prime}_{1}$ and $G^{\prime\prime}_{1}$) are inadequate in the non-linear regime and give only a measure of the average viscoelastic behaviourewoldt2008.
@@
Another way of probing non-linearity is through the elastic and viscous Lissajous (or Lissajous–Bowditch) curves, in which the raw stress is plotted as a function of the instantaneous strain and strain rate, respectively. A Hookean solid would return a positive diagonal elastic Lissajous curve and a perfectly circular viscous Lissajous curve, whereas a purely viscous material would yield the inverse. A linear viscoelastic response is then given by perfect ellipses in both elastic and viscous Lissajous, and distortions of the ellipse shape characterise a non-linear response, as higher harmonics affect the stress signal. cho2005 proposed decomposing the raw stress in LAOS tests into elastic and viscous components, which effectively collapse the Lissajous into one-dimensional curves, facilitating the analysis. Additionally, the Lissajous curves can also provide quantitative information. ewoldt2008 defined the minimum strain modulus as the tangent modulus at minimum strain, $G^{\prime}_{M}=d\sigma/d\gamma|_{\gamma=0}$, and the large strain modulus as the secant modulus at maximum strain, $G^{\prime}_{L}=\sigma/\gamma|_{\gamma=\gamma_{0}}$. Similarly, minimum and large rate dynamic viscosities can be defined as $\eta^{\prime}_{M}=d\sigma/d\dot{\gamma}|_{\dot{\gamma}=0}$ and $\eta^{\prime}_{L}=\sigma/\dot{\gamma}|_{\dot{\gamma}=\dot{\gamma}_{0}}$, respectively. As such, in the non-linear regime, these measures help quantify the (intracycle) non-linear behaviour, while in the linear regime, they reduce to $G^{\prime}_{1}$ and $\eta^{\prime}_{1}=G^{\prime\prime}_{1}/\omega$, respectively.
@@
ewoldt2008 also proposed to fit Chebyshev polynomials of the first kind to the decoupled stresses to allow an analysis of the harmonic contributions independently of the number of selected harmonics, which is usually arbitrary. The first order elastic and viscous Chebyshev coefficients are related to the widely used first-harmonic moduli: $e_{1}=G^{\prime}_{1}$ and $v_{1}=\eta^{\prime}_{1}=G^{\prime\prime}_{1}/\omega$, respectively, which are the sole contributors in the LVE; and the non-linear response can be described through the third order coefficients. A positive third-order elastic Chebyshev coefficient, $e_{3}>0$, relates to a higher elastic stress at maximum strain than the linear prediction (considering $e_{1}$ alone), whereas a negative value, $e_{3}<0$, corresponds to a lesser elastic stress at maximum strain than the linear representation. The same can be gathered from the third-order viscous coefficient regarding the viscous stress to the strain rate. This has led $e_{3}>0$ and $e_{3}<0$ to be respectively associated with intracycle elastic strain stiffening and softening, and $v_{3}>0$ and $v_{3}<0$, to intracycle viscous shear thickening and thinning, which often contradict the average viscoelastic behaviour described by $G^{\prime}_{1}$ and $G^{\prime\prime}_{1}$. This issue has been addressed beforemermet2015, ewoldt2013 and an alternative framework has been suggested that allows the interpretation of the third order coefficients within the context of the first-harmonic non-linear behaviourewoldt2013. Because the strain and strain rate are orthogonal by definition, the sign of the third order coefficients may be understood, not as a description of the non-linear response itself, but as an indicator of the deformation input that leads to the average non-linearity. For example, a positive $e_{3}$ relates to a relatively larger elastic stress at maximum strain ($\gamma=\gamma_{0}$) or minimum rate ($\dot{\gamma}=0$); as such, if the material presents elastic softening (that is, a decrease of $G^{\prime}_{1}$), $e_{3}>0$ indicates that the elastic softening is driven by the strain rate. The same can be said of the viscous response; a positive $v_{3}$ with an average viscous thinning material ($G^{\prime\prime}_{1}$ decrease) relates to the thinning non-linearity being driven by the strain, as the viscous stress is relatively lesser at $\gamma=\gamma_{0}$.
@@
FT-Chebyshev analysis is a powerful tool to unveil and quantify the intricacies of a material’s non-linear response. However, there is still a lack of in-depth analysis of non-linear magnetorheological data in the literature. As far as we know, the work of wang2021 is one of the few studies that probe experimental magnetorheological data through FT-Chebyshev analysis. The authors performed LAOS tests with a magnetorheological grease (a suspension of carbonyl iron particles in a semi-solid grease) under different magnetic field intensities. The authors found the application of the magnetic field transformed the rheological response from type I to type IIIhyun2002, and the LVE was significantly shortened, widening with magnetic field intensity increase but still narrower than for the non-magnetised grease. In the absence of magnetic influence, the MR grease is already highly viscoelastic, displaying elastic softening and viscous thinning through the decrease of $G^{\prime}_{1}$ and $\eta^{\prime}_{1}$, both driven by the strain rate ($G^{\prime}_{M}<G^{\prime}_{L}$ and $e_{3}>0$, $\eta^{\prime}_{L}<\eta^{\prime}_{M}$ and $v_{3}<0$). The magnetic field application led to significant alterations in the non-linear response. The authors report a sign shift of $e_{3}$ to negative values at larger strain (for intermediate fields) and positive $v_{3}$ at small strain, later shifting to $v_{3}<0$. The authors discuss the possibility of complex response arising from interactions between the fibre structure of the continuous phase and the magnetically-induced particle-chain structure.